OpenStudy (thefluffmuncher):

1. Solve by the linear combination method (with or without multiplication). 2x + 8y = –2 5x + 6y = 9 (Points : 1) (–1, 3) (3, –1) (–2, 2) (–2, 7)

OpenStudy (anonymous):

ill help

OpenStudy (thefluffmuncher):

thanks

OpenStudy (anonymous):

Solve for e for the first equation. e = -d + 2 Substitute that for the second equation and solve for d: d(-d + 2) = 4 -d² + 2d = 4 -d² + 2d - 4 = 0 By quadratic formula, which is d = (-b² ± √(b² - 4ac))/2a: a = -1 b = 2 c = -4 Substituting those values to the formula gives: d = (-2 ± √(2² - 4(-1)(-4)))/2(-1) d = (-2 ± √(4 - 16))/-2 d = (-2 ± √-12)/-2 Since the discriminant b² - 4ac gives nonreal solutions, there are no real solutions for the system of equations. 2. y = 3x - 7 6x - 2y = 12 Substitute y with 3x - 7 for the second equation: 6x - 2(3x - 7) = 12 6x - 6x + 14 = 12 0x + 14 = 12 0x = -2 x = undefined Since 12 is not divisible by 0, x is not defined. Hence, there is also no solution for that system of equations. 3. 3x + 2y = 11 x - 2 = -4y Solve for x for the second equation: x = 2 - 4y Substitute that for the first equation: 3(2 - 4y) + 2y = 11 6 - 12y + 2y = 11 6 - 10y = 11 -10y = 11 - 6 -10y = 5 y = -½ Substitute that for either equation: x - 2 = -4(-½) x - 2 = 2 x = 2 + 2 x = 4 Hence, the solution is (4,-½). I hope this helps!

OpenStudy (thefluffmuncher):

i haven't even reached quadratic equation yet

OpenStudy (campbell_st):

The linear combination method is basically the old elimination method multiply each of the equations so that 1 term can be eliminated then add the 2 equations. looking at the problem I would multiply the 1st equation by 5 and the 2nd equation by -2 10x + 40y = -10 -10x -12y = -18 ----------------- add the equations and x will be eliminated you can do this bit then solve for y. when you get y substitute into either of the original equations to find x.

OpenStudy (thefluffmuncher):

i was wondering of this correct?